The water level of a pond drops an inch every 7 days without rain. How many days will it take the pond's water level to drop by 12 inches?
Answer by TimothyLamb(4379)

---
h/d = 1/7
---
12/d = 1/7
d = 12*7
d = 84 days
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php